OpenStudy (anonymous):

You have 13 grams of chlorine. How many grams of iron (III) chlorine are produced

OpenStudy (anonymous):

1st find the moles of chlorine gas. i.e 13/71=0.183 2Fe + 3Cl2 ----> 2FeCl3 Now. 3moles of Cl2 form = 2 moles of FeCl3 0.183 moles of Cl2 can form = 2/3 * 0.183 = 0.122 moles of Fecl3. Mass of FeCl3 = moles * molar mass = 0.122 * 55.8 = 6.8 g of FeCl3.